OpenStudy (anonymous):

Given f(x) = 5x - 6 and g(x) = x2 - 5x + 6, find the function (f + g)(x). Use the ^ key for the exponent. Enter your answer as the example: 4x^3.

OpenStudy (hoblos):

(f + g)(x)= 5x - 6 + x^2 - 5x + 6 = x^2

OpenStudy (anonymous):

thank you ^.^

OpenStudy (hoblos):

my pleasure XD

OpenStudy (anonymous):

the 5 and 6 just cancel out?

OpenStudy (hoblos):

yes

OpenStudy (anonymous):

okay thank you^.^

OpenStudy (hoblos):

:D